# Improved Generation of Adversarial Examples Against Safety-aligned LLMs

### v1

@@ Abstract
The remarkable performance of large language models (LLMs) has raised both interest and concern regarding their safety and trustworthiness. Despite numerous efforts to ensure they adhere to safety standards and produce harmless content, some successes have been achieved in bypassing these restrictions, known as jailbreak attacks against LLMs. Adversarial prompts (or say, adversarial examples) generated using gradient-based methods exhibit outstanding performance in performing jailbreak attacks automatically. Nevertheless, due to the discrete nature of texts, the input gradient of LLMs struggles to precisely reflect the magnitude of loss change that results from token replacements in the prompt, leading to limited attack success rates against safety-aligned LLMs, even in the white-box setting. In this paper, we explore a new perspective on this problem, suggesting that it can be alleviated by leveraging innovations inspired in transfer-based attacks that were originally proposed for attacking black-box image classification models. For the first time, we appropriate the ideologies of effective methods among these transfer-based attacks, i.e., Skip Gradient Method [48] and Intermediate Level Attack [16], for improving the effectiveness of automatically generated adversarial examples against white-box LLMs. With appropriate adaptations, we inject these ideologies into gradient-based adversarial prompt generation processes and achieve significant performance gains without introducing obvious computational cost. Meanwhile, by discussing mechanisms behind the gains, new insights are drawn, and proper combinations of these methods are also developed. Our empirical results show that $87\%$ of the query-specific adversarial suffixes generated by the developed combination can induce Llama-2-7B-Chat to produce the output that exactly matches the target string on AdvBench. This match rate is $33\%$ higher than that of a very strong baseline known as GCG, demonstrating advanced discrete optimization for adversarial prompt generation against LLMs. In addition, without introducing obvious cost, the combination achieves $>30\%$ absolute increase in attack success rates compared with GCG when generating both query-specific ($38\%\rightarrow 68\%$) and universal adversarial prompts ($26.68\%\rightarrow 60.32\%$) for attacking the Llama-2-7B-Chat model on AdvBench. Code at: https://github.com/qizhangli/Gradient-based-Jailbreak-Attacks.
@@ 1 Introduction
One of the main difficulties in generating adversarial examples for NLP models lies in the fact that text is discrete by nature, making it challenging to use gradient-based optimization methods to devise adversarial attacks. There has been some work [13, 47, 38, 18, 52] attempted to overcome this issue. For instance, recently, a method called Greedy Coordinate Gradient (GCG) attack [52] has shown significant jailbreaking improvements, by calculating gradients of cross-entropy loss w.r.t. one-hot representations of chosen tokens in a prompt and utilizing them in a greedy manner. However, due to the fact that the gradients w.r.t. one-hot vectors do not provide precise indication of the loss change that results from a token replacement, the GCG attack shows limited white-box attack success rates against some safty-aligned LLMs, e.g., Llama-2-Chat models [41].
In this paper, we carefully examine the discrepancy between the gradient of the adversarial loss w.r.t. one-hot vectors and the real effect of the change in loss that results from token replacement. We present a new perspective that this gap resembles the gap between input gradients calculated using a substitute model and the real effect of perturbing inputs on the prediction of a black-box victim model, which has been widely studied in transfer-based attacks against black-box image classification models [40, 34, 26, 16, 48, 14, 24]. Based on this new perspective, for the first time, we attempt to appropriating the ideologies of two effective methods among these transfer-based methods, i.e., Skip Gradient Method (SGM) [48] and Intermediate Level Attack (ILA) [16], to improve the gradient-based attacks against LLMs. With appropriate adaptations, we successfully inject these ideologies into the gradient-based adversarial prompt generation without additional computational cost. By discussing the mechanisms behind the advanced performance, we provide new insights about improving discrete optimizations on LLMs. Moreover, we provide an appropriate combination of these methods. The experimental results demonstrate that $87\%$ of the query-specific adversarial suffixes generated by the combination for attacking Llama-2-7B-Chat on AdvBench can induce the model output exact target string, which outperforms a strong baseline named GCG attack (54%), indicating an advanced discrete optimization for adversarial prompt generation against LLMs. In addition, the combination achieves attack success rates of 68% for query-specific and 60.32% for universal adversarial prompt generation when attacking Llama-2-7B-Chat on AdvBench, which are higher than those of the baseline GCG (38% and 26.68%).
@@ 2 Related Work
Transfer-based attacks. Transfer-based attacks attempt to craft adversarial examples on a white-box substitute model to attack the black-box victim model, by leveraging the transferability of adversarial examples [40], which is a phenomenon that adversarial examples crafted on a white-box substitute model can also mislead the unknown victim models with a decent success rate. The transfer-based attacks have been thoroughly investigated in the setting of attacking image classification models [20, 49, 6, 48, 16, 11, 14, 15, 23, 24]. Some recent methods also utilize the transferability of adversarial prompts to perform black-box attacks against LLMs [52, 25, 39]. While in this work, we mainly focus on improving the white-box attack success rate. In our work, we reveal a closely relationship between the optimization in transfer-based attacks and discrete optimization of the gradient-based jailbreak attacks against LLMs. We then appropriate the ideologies of two effective transfer-based attack methods developed in the setting of attacking image classification model, i.e., SGM [48] and ILA [16]. Moreover, by adapting these strategies and analyzing the mechanism behind them, we provide some new insights about potential solutions for addressing problems involving discrete optimization in NLP models with transformer architecture, e.g., prompt tuning.
@@ 3.1 Rethinking Gradient-based Attacks Against LLMs
A typical LLM $f:\mathcal{X}\rightarrow\mathbb{R}^{|V|}$ with a vocabulary $V$ is trained to map a sequence of tokens $x_{1:n}=[x_{1},...,x_{n}]\in\mathcal{X},x_{i}\in V$ to a probability distribution over the next token, denoted as $p_{f}(x_{n+1}|x_{1:n})$. To evoke a jailbreak to induce the a safety-aligned LLM generate harmful content according the user query, GCG attempts to add an adversarial suffix to the original user query and iteratively modifies the adversarial suffix to encourages the model output an affirmative target phrase, e.g., “Sure, here’s …”. Consider an adversarial prompt (which is the concatenation of a user query and an adversarial suffix) as $x_{1:n}$ and a further concatenation with a target phrase as $x_{1:n^{*}}$, GCG aims to minimize the adversarial loss, which corresponds to the negative log probability of the target phrase. It can be written as
$\min_{x_{\mathcal{A}}\in\{1,\dots,V\}^{|\mathcal{A}|}}\frac{1}{n^{*}-n}\sum_{i=1}^{n^{*}-n}-\log p_{f}(x_{n+i}|x_{1:n+i-1}),$ (1)
@@
where $x_{\mathcal{A}}$ denotes the tokens of adversarial suffix in $x_{1:n}$ and $\mathcal{A}$ denotes the set of indices of the adversarial suffix tokens, and we use $L(x)$ to denote the adversarial loss. At each iteration, it computes the gradient of the cross-entropy loss w.r.t. the one-hot vector of a single token and uses each value of gradient to estimate the effect of replacing the current token with the corresponding one on loss. Since the discrete nature of input, there is a gap between input gradients and the effects of token replacements, thus the estimate is not accurate enough. Previous efforts attempt to solve this problem by collecting a candidate set of token replacements according to the Top-$k$ values in the gradient and evaluating the candidates to pick the best token replacement with minimal loss [52, 38]. We target this problem in another fashion by refining the input gradient to narrow the gap.
@@
We attempt to solve this problem from a new perspective that treats the estimation as using a “surrogate model” which takes continuous inputs to attack the “real model” which takes discrete text inputs. Let us provide explanations as follows. The strategy of utilizing the gradient w.r.t. the one-hot vector is actually treating the one-hot vector as though it was a continuous variable. Thus, we consider that the gradient-based attacks essentially build a surrogate model, $f^{S}:\mathbb{R}^{n\times|V|}\rightarrow\mathbb{R}^{|V|}$, which has the same architecture and parameters as $f$, except it maps input from a continuous space to the probability distribution. This is different from the real model $f$, which maps its input from a discrete space to the probability distribution. Then, the gradient-based attacks use the input gradient computed on the “surrogate model” to estimate the effect of the loss change resulting from token replacements on the “real model”. Therefore, the problem of inaccurate estimates is similar to the challenges encountered in transfer-based attacks on image classification models. Specifically, in the context of transfer-based attacks on image classification models, the goal is to compromise an unknown victim model. To achieve this, the attackers employ a substitute model that performs the same task as the victim model, and use this substitute model to calculate the gradient of the classification loss w.r.t. the input. Such input gradient estimates how changes in the input affect the changes in the prediction on the victim model. The attackers then modify the adversarial example to increase the classification loss on the victim model. Many methods [7, 49, 16, 48, 14, 23] have been developed to refine the gradient computed on the substitute model, in order to narrow the gap between the input gradients and the real effects of loss change results from changes in the inputs on the victim model. This new perspective allows one to introduce a series of innovations developed within the realm of transfer-based attacks on image classification models to refine the gradient computation in the context of gradient-based adversarial prompt generation against safety-aligned LLMs.
@@ 3.2 Reducing the Gradients from Residual Modules
An $l$-layers LLM can be decomposed into $2l$ residual building blocks, with each block consisting of a residual module (which should be an MLP or an attention module) and a parallel skip connection branch as illustrated in Figure 2. The $m$-th block maps an intermediate representation $z_{m}$ to $z_{m+1}$, i.e., $z_{m+1}=I(z_{m})+R_{m}(z_{m})$, where $I$ is an identity function representing the skip connection branch and $R_{m}$ denotes the residual module of the $m$-th block. By adopting a decay factor $\gamma\in[0,1]$ for the residual modules, SGM calculates the derivative of the $m$-th block as $\nabla_{z_{m}}z_{m+1}=1+\gamma\nabla_{z_{m}}R_{m}(z_{m})$. We incorporate this strategy into gradient-based automatic adversarial prompt generation, denoted as Language SGM (LSGM). We evaluate the performance of integrating this strategy into GCG and AutoPrompt attacks by setting $\gamma=0.5$, and show the results in Figure 2. The experiment is conducted to generate query-specific adversarial suffixes against Llama-2-7B-Chat [41] on the first 100 harmful queries in AdvBench [52]. To provide a more comprehensive comparison, we report not only the adversarial loss but also the fraction of adversarial prompts that result in outputs exactly matching the target string, dubbed match rate, which is also used as an evaluation metric in the paper of GCG [52]. It can be seen from the figure that reducing gradients from residual modules can indeed improve the performance of gradient-based adversarial prompt generation. The GCG-LSGM achieves a match rate of $72\%$, while the baseline (i.e., GCG) only obtains $54\%$ match rate. We also evaluate the attack success rate (ASR) by using the evaluator proposed by HarmBench [27]. It shows an ASR of $62\%$ when using GCG-LSGM, while the GCG only achieves $38\%$. The results confirm that reducing gradients from residual modules helps the gradients w.r.t. one-hot representations to be more effective to indicate the real effects of token replacements.
@@ 3.3 Adapting Intermediate Level Attack for Gradient-base Attacks Against LLMs
Let $h_{r}$ be the intermediate representation at the $r$-th layer, and the directional guide $v_{r}=\hat{h}^{t}_{r}-\hat{h}^{0}_{r}$ is obtained by a $t$-iterations preliminary attack (e.g., GCG). To directly apply ILA, we maximize the scalar projection of the intermediate representation $h_{r}$ onto the directional guide, i.e., maximize $L_{\mathrm{ILA}}(x)=v_{r}^{T}h_{r}$. We obtain the directional guide from the GCG attack and then evaluate the direct application of ILA using the GCG attack as the baseline. We find that it exhibits deteriorated performance compared with GCG. Specifically, it achieves a match rate of $44\%$, while the original GCG obtains $54\%$ match rate. This result demonstrates that the ILA fails to be directly applied to gradient-based adversarial prompt generation.
(a) $h_{r}$ and $h_{r,n}$
(b) $h_{r,o}$
Figure 5: Negative average PCCs (with higher values indicating stronger negative correlations) computed on (a) the entire intermediate representation $h_{r}$ and the last adversarial prompt token representation $h_{r,n}$, and (b) the token representations $h_{r,o}$, at the $r$-th layer. Best viewed in color.
@@
Now, let us discuss the reason why directly introducing ILA fails to improve the performance. Recall that the intermediate level attacks against image classification models essentially assume that the scalar projection of intermediate representation onto the directional guide has a positive correlation with the transferability of adversarial examples, which means that the larger the scalar projection obtained, the more transferability (higher classification loss on victim models) the adversarial example achieved [22]. Back to our setting, direct application of ILA requires that there is a negative correlation between the scalar projection of the intermediate representation onto the directional guide and the adversarial loss (since we aim to minimize the adversarial loss in contrast to the setting of non-target attack against image classification model). We conduct experiments to verify whether this assumption holds in attacking LLMs. We use Pearson’s correlation coefficient (PCC) to show the correlation between them. With a range in $[-1,1]$, PCC close to $1$ indicates a positive correlation, and PCC close to $-1$ means a negative correlation. The experiment is performed with following steps. Firstly, we perform GCG attack to obtain an adversarial example and a directional guide $v_{r}$. Next, we randomly alter adversarial tokens several times and pass these new prompts into the model to collect a set of scalar projections paired with their corresponding adversarial losses, denoted as $\{(v_{r}^{T}\tilde{h}^{c^{*}}_{r},L(\tilde{x}^{c^{*}})\}$. Then we can calculate a PCC between the scalar projections and loss values on this set. By doing this experiment on different examples, we can obtain the average PCC. In Figure 5(a), we show the negative average PCCs (with higher values indicating stronger negative correlations) between the scalar projection and the adversarial loss at different layers. The negative average PCCs are only between 0.4 and 0.5, showing low correlation between them. It confirms that the failure to apply ILA directly to gradient-based attacks against LLMs is due to the low correlation between the scalar projection of the intermediate representation onto the directional guide and the adversarial loss. This phenomenon is distinct from the one observed in the setting of attacking image classification models [22].
@@
Figure 6: An example of the internal computation of an LLM. The $\ell_{o}$ represents cross-entropy loss.
@@
Nevertheless, by unfolding the internal computation of an LLM as a grid of token representations $\{h_{r,o}\}$, where $h_{r,o}$ denotes the $o$-th token representation at $r$-th layer, as depicted in Figure 7, some previous work [29, 30, 9, 32] have shown that different token representations at a middle layer show distinct effects on the model output. It inspires us to investigate whether the negative correlation exists between the scalar projection of a single token representation $h_{r,o}$ onto the corresponding directional guide $v_{r,o}$ and the adversarial loss. Note that during the optimization, the input is a concatenation of the adversarial prompt tokens $x_{1:n}$ and the target tokens $x_{n+1:n^{*}}$, and the adversarial loss is the average over $n^{*}-n$ cross-entropy losses. Consider $L(x)=1/(n^{*}-n)\sum_{o=n}^{n^{*}-1}\ell_{o}$, where $\ell_{o}=-\log p_{f}(x_{o+1}|x_{1:o})$, for a token representation of $h_{r,o^{*}},o^{*}\in\{n,\dots,n^{*}-1\}$, it only affects the loss values of $\{\ell_{o^{*}},\dots,\ell_{n^{*}-1}\}$, hence we only average these loss values as the corresponding adversarial loss to evaluate the correlation. Since the numbers of target tokens are different across different adversarial prompts, we select the first 10 tokens from the target tokens to compute the PCCs. In Figure 5(b), we show the negative average PCCs corresponding to single token representations $\{h_{r,n-|\mathcal{A}|+1},\dots,h_{r,n^{*}-1}\}$ at different middle layers. Interestingly, we can see a high negative correlation between the scalar projection of the last adversarial prompt token representation onto the directional guide, i.e., $v_{r,n}^{T}h_{r,n}$, and the adversarial loss, i.e., $L(x)$, at some middle layers. To clearly demonstrate the difference with the scalar projection of the entire representation onto the directional guide and the adversarial loss, we compare them in Figure 5(a). It can be seen that from the 8-th layer onward, the negative average PCCs of $h_{r,n}$ become larger than those of $h_{r}$ and then exceed $0.7$ starting from the 16-th layer, showing a high negative correlation. We conjecture that this phenomenon is attributed to two factors. Firstly, the last adversarial prompt token representation contains information from the entire adversarial prompt, making it more representative than the preceding token representations. Secondly, since the last adversarial prompt token representation corresponds to the generation of the first target token, which is crucial for the generation of subsequent target tokens [2], the changes in the losses of subsequent target tokens are primarily affected by the changes in the last adversarial prompt token representation.

### v2

@@ Abstract
Adversarial prompts (or say, adversarial examples) generated using gradient-based methods exhibit outstanding performance in performing automatic jailbreak attacks against safety-aligned LLMs. Nevertheless, due to the discrete nature of texts, the input gradient of LLMs struggles to precisely reflect the magnitude of loss change that results from token replacements in the prompt, leading to limited attack success rates against safety-aligned LLMs, even in the white-box setting. In this paper, we explore a new perspective on this problem, suggesting that it can be alleviated by leveraging innovations inspired in transfer-based attacks that were originally proposed for attacking black-box image classification models. For the first time, we appropriate the ideologies of effective methods among these transfer-based attacks, i.e., Skip Gradient Method [53] and Intermediate Level Attack [18], into gradient-based adversarial prompt generation and achieve significant performance gains without introducing obvious computational cost. Meanwhile, by discussing mechanisms behind the gains, new insights are drawn, and proper combinations of these methods are also developed. Our empirical results show that $87\%$ of the query-specific adversarial suffixes generated by the developed combination can induce Llama-2-7B-Chat to produce the output that exactly matches the target string on AdvBench. This match rate is $33\%$ higher than that of a very strong baseline known as GCG, demonstrating advanced discrete optimization for adversarial prompt generation against LLMs. In addition, without introducing obvious cost, the combination achieves $>30\%$ absolute increase in attack success rates compared with GCG when generating both query-specific ($38\%\rightarrow 68\%$) and universal adversarial prompts ($26.68\%\rightarrow 60.32\%$) for attacking the Llama-2-7B-Chat model on AdvBench. Code at: https://github.com/qizhangli/Gradient-based-Jailbreak-Attacks.
@@ 1 Introduction
One of the main difficulties in generating adversarial examples for NLP models lies in the fact that text is discrete by nature, making it challenging to use gradient-based optimization methods to devise adversarial attacks. There has been some work [15, 52, 43, 20, 58] attempted to overcome this issue. For instance, recently, a method called Greedy Coordinate Gradient (GCG) attack [58] has shown significant jailbreaking improvements, by calculating gradients of cross-entropy loss w.r.t. one-hot representations of chosen tokens in a prompt and replacing them in a greedy manner. However, due to the fact that the gradients w.r.t. one-hot vectors do not provide precise indication of the loss change that results from a token replacement, the GCG attack shows limited white-box attack success rates against some safty-aligned LLMs, e.g., Llama-2-Chat models [46].
In this paper, we carefully examine the discrepancy between the gradient of the adversarial loss w.r.t. one-hot vectors and the real effect of the change in loss that results from token replacement. We present a new perspective that this gap resembles the gap between input gradients calculated using a substitute model and the real effect of perturbing inputs on the prediction of a black-box victim model, which has been widely studied in transfer-based attacks against black-box image classification models [45, 39, 31, 18, 53, 16, 29, 28]. Based on this new perspective, for the first time, we attempt to appropriating the ideologies of two effective methods among these transfer-based methods, i.e., Skip Gradient Method (SGM) [53] and Intermediate Level Attack (ILA) [18], to improve the gradient-based attacks against LLMs. With appropriate adaptations, we successfully inject these ideologies into the gradient-based adversarial prompt generation without additional computational cost. By discussing the mechanisms behind the advanced performance, we provide new insights about improving discrete optimizations on LLMs. Moreover, we provide an appropriate combination of these methods. The experimental results demonstrate that $87\%$ of the query-specific adversarial suffixes generated by the combination for attacking Llama-2-7B-Chat on AdvBench can induce the model output exact target string, which outperforms a strong baseline named GCG attack (54%), indicating an advanced discrete optimization for adversarial prompt generation against LLMs. In addition, the combination achieves attack success rates of 68% for query-specific and 60.32% for universal adversarial prompt generation when attacking Llama-2-7B-Chat on AdvBench, which are higher than those of the baseline GCG (38% and 26.68%).
@@ 2 Related Work
Transfer-based attacks. Transfer-based attacks attempt to craft adversarial examples on a white-box substitute model to attack the black-box victim model, by leveraging the transferability of adversarial examples [45], which is a phenomenon that adversarial examples crafted on a white-box substitute model can also mislead the unknown victim models with a decent success rate. The transfer-based attacks have been thoroughly investigated in the setting of attacking image classification models [22, 54, 8, 53, 18, 13, 16, 17, 27, 26, 25, 28]. Some recent methods also utilize the transferability of adversarial prompts to perform black-box attacks against LLMs [58, 30, 44]. While in this work, we mainly focus on improving the white-box attack success rate. In our work, we reveal a closely relationship between the optimization in transfer-based attacks and discrete optimization of the gradient-based jailbreak attacks against LLMs. We then appropriate the ideologies of two effective transfer-based attack methods developed in the setting of attacking image classification model, i.e., SGM [53] and ILA [18]. Moreover, by adapting these strategies and analyzing the mechanism behind them, we provide some new insights about potential solutions for addressing problems involving discrete optimization in NLP models with transformer architecture, e.g., prompt tuning.
@@ 3.1 Rethinking Gradient-based Attacks Against LLMs
A typical LLM $f:\mathcal{X}\rightarrow\mathbb{R}^{|V|}$ with a vocabulary $V$ is trained to map a sequence of tokens $x_{1:n}=[x_{1},...,x_{n}]\in\mathcal{X},x_{i}\in V$ to a probability distribution over the next token, denoted as $p_{f}(x_{n+1}|x_{1:n})$. To evoke a jailbreak to induce the a safety-aligned LLM generate harmful content according the user query, GCG attempts to add an adversarial suffix to the original user query and iteratively modifies the adversarial suffix to encourages the model output an affirmative target phrase, e.g., “Sure, here’s …”. Consider an adversarial prompt (which is the concatenation of a user query and an adversarial suffix) as $x_{1:n}$ and a further concatenation with a target phrase as $x_{1:n^{*}}$, GCG aims to minimize the adversarial loss $L(x_{1:n^{*}})$ (denoted as $L(x)$ for simplicity), which corresponds to the negative log probability of the target phrase. It can be written as
$\min_{x_{\mathcal{A}}\in\{1,\dots,V\}^{|\mathcal{A}|}}L(x_{1:n^{*}})=\min_{x_{\mathcal{A}}\in\{1,\dots,V\}^{|\mathcal{A}|}}\frac{1}{n^{*}-n}\sum_{i=1}^{n^{*}-n}-\log p_{f}(x_{n+i}|x_{1:n+i-1}),$ (1)
@@
where $x_{\mathcal{A}}$ denotes the tokens of adversarial suffix in $x_{1:n}$ and $\mathcal{A}$ denotes the set of indices of the adversarial suffix tokens. At each iteration, it computes the gradient of the adversarial loss w.r.t. the one-hot vector of a single token and uses each value of gradient to estimate the effect of replacing the current token with the corresponding one on loss. Since the discrete nature of input, there is a gap between input gradients and the effects of token replacements, thus the estimate is not accurate enough. Previous efforts attempt to solve this problem by collecting a candidate set of token replacements according to the Top-$k$ values in the gradient and evaluating the candidates to pick the best token replacement with minimal loss [58, 43]. Due to the large gap, they requires a large candidate set size, e.g., 512 in GCG, which result in a large computational cost. Ideally, if the input gradients can accurately reflect the effects of token replacements, $k=1$ is sufficient to achieve minimal loss without needing a candidate set, thus obtaining the optimal token replacement with the lowest computational cost. Therefore, we advocate for refining the input gradient to narrow the gap, thereby improving performance while reducing computational cost.
@@
We attempt to introduce a new perspective to the gap between the input gradients and the real effects of token replacements. Let us first revisit transfer-based attacks on image classification models. To generate an adversarial example that misleads an unknown victim model, where the input gradient is not accessible, attackers use a white-box substitute model as a proxy for the victim model and utilize its input gradient. Due to the gap between the input gradient of the substitute model and the real input gradient of the victim model, directly using the input gradient of the substitute model to modify the example yields unsatisfactory results. Efforts [9, 54, 18, 53, 24, 16, 27] have been made to refine the gradient computed on the substitute model in order to narrow this gap. Returning to our discrete optimization problem in the adversarial prompt generation, the strategy of utilizing the gradient w.r.t. the one-hot vector of the token is actually treating the one-hot vector as though it was a continuous variable, resulting in the gap between this gradient and the “real gradient” (e.g., the real effects of token replacements). We consider that this gap is analogous to the input gradient gap between substitute and victim models in the context of transfer-based attacks. This new perspective allows one to introduce a series of innovations developed within the realm of transfer-based attacks on image classification models to refine the gradient computation in the context of gradient-based adversarial prompt generation against safety-aligned LLMs.
@@ 3.2 Reducing the Gradients from Residual Modules
An $l$-layers LLM can be decomposed into $2l$ residual building blocks, with each block consisting of a residual module (which should be an MLP or an attention module) and a parallel skip connection branch as illustrated in Figure 2. The $m$-th block maps an intermediate representation $z_{m}$ to $z_{m+1}$, i.e., $z_{m+1}=I(z_{m})+R_{m}(z_{m})$, where $I$ is an identity function representing the skip connection branch and $R_{m}$ denotes the residual module of the $m$-th block. By adopting a decay factor $\gamma\in[0,1]$ for the residual modules, SGM calculates the derivative of the $m$-th block as $\nabla_{z_{m}}z_{m+1}=1+\gamma\nabla_{z_{m}}R_{m}(z_{m})$. We incorporate this strategy into gradient-based automatic adversarial prompt generation, denoted as Language SGM (LSGM). We evaluate the performance of integrating this strategy into GCG and AutoPrompt attacks by setting $\gamma=0.5$, and show the results in Figure 2. The experiment is conducted to generate query-specific adversarial suffixes against Llama-2-7B-Chat [46] on the first 100 harmful queries in AdvBench [58]. To provide a more comprehensive comparison, we report not only the adversarial loss but also the fraction of adversarial prompts that result in outputs exactly matching the target string, dubbed match rate, which is also used as an evaluation metric in the paper of GCG [58]. It can be seen from the figure that reducing gradients from residual modules can indeed improve the performance of gradient-based adversarial prompt generation. The GCG-LSGM achieves a match rate of $72\%$, while the baseline (i.e., GCG) only obtains $54\%$ match rate. We also evaluate the attack success rate (ASR) by using the evaluator proposed by HarmBench [32]. It shows an ASR of $62\%$ when using GCG-LSGM, while the GCG only achieves $38\%$. On the other hand, when reducing the gradient from the skip connection, both the match rate and the attack success rate will drop to 0%. The results confirm that reducing gradients from residual modules helps the gradients w.r.t. one-hot representations to be more effective to indicate the real effects of token replacements.
@@ 3.3 Adapting Intermediate Level Attack for Gradient-base Attacks Against LLMs
Let $h_{r}$ be the intermediate representation at the $r$-th layer, and the directional guide $v_{r}=\hat{h}^{0}_{r}-\hat{h}^{t}_{r}$ is obtained by a $t$-iterations preliminary attack (e.g., GCG). Note that since our objective is to minimize the adversarial loss instead of maximizing the victim’s classification loss, as in the original setting of ILA paper [18], the definition of the directional guide is the opposite of that in the ILA paper. According the implementation of ILA, we maximize the scalar projection of the intermediate representation $h_{r}$ onto the directional guide $v_{r}$, i.e., maximize $L_{\mathrm{ILA}}(x)=h^{T}_{r}v_{r}$. However, we evaluated ILA using GCG as the back-end method and found that it exhibited deteriorated performance compared to the baseline. Specifically, when attacking Llama-2-7B-Chat, it achieves a match rate of $44\%$, while the baseline GCG obtains $54\%$ match rate. This result demonstrates that when applied directly, the ILA fails to facilitate gradient-based adversarial prompt generation.
(a) PCCs computed on the $h_{r}$
(b) PCCs computed on the $h_{r,o}$
Figure 5: (a) The PCCs computed on the entire intermediate representations, i.e., $\mathrm{PCC}(h^{T}_{r}\nabla_{h_{r}}L(x),L(x))$ and $\mathrm{PCC}(h^{T}_{r}v_{r},L(x))$, at different layers of Llama-2-7B-Chat. (b) The PCCs computed on the $o$-th token intermediate representations, i.e., $\mathrm{PCC}(h_{r,o}\nabla_{h_{r,o}}L(x),L(x))$ and $\mathrm{PCC}(h^{T}_{r,o}v_{r,o},L(x))$, at the mid-layer of Llama-2-7B-Chat. Best viewed in color.
@@
Let us discuss why directly introducing ILA fails to improve performance. Recall that intermediate-level attacks against image classification models essentially assume that the scalar projection of an intermediate representation onto the directional guide has a positive correlation with the transferability of adversarial examples, which means that the larger the scalar projection obtained, the greater the transferability (i.e., higher classification loss on victim models) the adversarial example achieves [26]. Previous work [26] has also shown that the ILA is equivalent to replacing the gradient w.r.t. the intermediate representation with the directional guide. This demonstrates that ILA necessitates a stronger positive correlation between the scalar projection of the representation onto the directional guide and the victim’s classification loss, compared to the positive correlation between the scalar projection of the representation onto its gradient and the victim’s classification loss. Back to our setting, we conduct experiments to examine whether these assumptions hold when attacking LLMs. We use Pearson’s correlation coefficient (PCC) to show the correlation between the scalar projection and adversarial loss. With a range in $[-1,1]$, a PCC close to $1$ indicates a positive correlation, while a PCC close to $-1$ means a negative correlation. The experiment is conducted with following steps. Firstly, we perform a GCG attack to obtain an adversarial example, and use it to derive a directional guide $v_{r}$ and the gradient of adversarial loss $L(x)$ w.r.t. the intermediate representation $h_{r}$, i.e.$\nabla_{h_{r}}L(x)$. Next, we randomly alter adversarial tokens several times and input these new prompts into the model to collect a set of intermediate representations paired with their corresponding adversarial losses. We then calculate the PCC between the scalar projection of the intermediate representation onto the directional guide and the adversarial loss, i.e., $\mathrm{PCC}(h^{T}_{r}v_{r},L(x))$, and the PCC between the scalar projection of the intermediate representation onto its gradient and the adversarial loss, i.e., $\mathrm{PCC}(h^{T}_{r}\nabla_{h_{r}}L(x),L(x))$. We perform this experiment using Llama-2-7B-Chat on AdvBench. In Figure 5(a), we show the PCCs between the scalar projection and the adversarial loss at different layers. It can be seen that in every layer, the PCCs computed using the directional guide are all $\leq 0.5$, showing a weaker correlation compared to the PCCs computed using the intermediate gradient, which exceed 0.8 after the 14th layer. This confirms that the failure to directly apply ILA is due to its poorer performance in the intermediate layers in indicating the change of adversarial loss compared to the baseline. This phenomenon is distinct from the one observed in the setting of attacking image classification models [26].
@@
Figure 6: An example of the internal computation of an LLM. The $\ell$ represents cross-entropy loss.
@@
Nevertheless, by unfolding the internal computation of an LLM as a grid of token representations $\{h_{r,o}\}$, where $h_{r,o}$ denotes the $o$-th token representation at $r$-th layer, as depicted in Figure 7, some previous work [34, 35, 11, 37] have shown that different token representations at a intermediate layer show distinct effects on the model output. It inspires us to further investigate whether the positive correlation exists between the scalar projection of a single token representation onto its directional guide and the adversarial loss (i.e., $\mathrm{PCC}(h^{T}_{r,o}v_{r,o},L(x))$), and whether this correlation is stronger than the positive correlation between the scalar projection computed on its gradient and the adversarial loss (i.e., $\mathrm{PCC}(h^{T}_{r,o}\nabla_{h_{r,o}}L(x),L(x))$). In Figure 5(a), we present $\mathrm{PCC}(h^{T}_{r,o}v_{r,o},L(x))$ and $\mathrm{PCC}(h^{T}_{r,o}\nabla_{h_{r,o}}L(x),L(x))$ with different choices of $o$-th token at the mid-layer of Llama-2-7B-Chat. Since the number of target tokens varies across different adversarial prompts, we select the first five tokens from the target phrase. First, the scalar projections computed using the directional guide and intermediate gradient both show a very weak positive correlation with the adversarial loss in the intermediate representations of adversarial tokens. Second, the PCCs calculated for the token representation of “]”, which is the last token of the input prompt and is expected to output the first token of the target phrase (e.g., “Sure”), demonstrate that the scalar projection computed using the gradient can better indicate the change in adversarial loss than that computed using the directional guide. We consider that successfully generating the first token of the target (e.g., “Sure”) plays a crucial role in reducing the adversarial loss. The adversarial prompt used to obtain the directional guide is not sufficiently optimized to minimize the loss associated with generating the first token of the target. Somewhat interestingly, the high PCC value between the projection of last token representation onto the gradient (or directional guide) and the adversarial loss indicates that directly moving the last token representation along the direction of gradient (i.e., intervening in the last token representation by adding a vector in the direction of gradient or directional guide) will effectively reduce the adversarial loss. Some concurrent work [2, 55] also observes that certain directions can be added to the token representations to jailbreak safety-aligned LLMs. In contrast to our approach, they collect the token representations of a set of harmful queries and a set of harmless queries, defining the direction as the difference between the average last token representations of these two sets. Third, for other token representations, the scalar projections onto directional guides show a stronger correlation with adversarial loss than the scalar projections onto their gradients. This indicates that projecting the token representation onto the directional guide can better reflect changes in adversarial loss compared to projecting it onto the intermediate gradient.
Therefore, to effectively adapt ILA for adversarial prompt generation we propose replacing the intermediate gradient of the adversarial loss with a directional guide at specific token representations, rather than at the entire intermediate representation as in the original implementation. We re-normalize the directional guide by $\|\nabla_{h_{r,o}}L(x)\|_{2}/\|v_{r,o}\|_{2}$, in order to avoid too large (or small) compared to the gradient w.r.t. other token representations. We denote this adaptation of ILA as Language ILA (LILA). When performing LILA at the $o$-th token representation at $r$-th layer, during backpropagation, the gradient w.r.t. the token representation, i.e., $\nabla_{h_{r}}L(x)$ is changed to
$\nabla^{\mathrm{LILA}}_{h_{r}}L(x)=[\nabla_{h_{r,1}}L(x),\nabla_{h_{r,2}}L(x),...,\frac{\|\nabla_{h_{r,o}}L(x)\|_{2}}{\|v_{r,o}\|_{2}}v_{r,o},...,\nabla_{h_{r,n^{*}-1}}L(x)].$ (2)